A 24.93 g sample of ice is heated from -16 oC to its melting point of 0 oC and then melted completely. Next it is raised to its boiling point of 100 oC and boiled completely. Finally the sample is heated to 113 oC. How much energy was absorbed by the sample? (Answer in kJ pay close attention to units!)

Sice=2.1 J/goC Delta Hf = 6.02 kJ/mol

Ssteam=1.7 J/goC Delta Hv = 2.26 kJ/g


1
Expert's answer
2021-03-25T04:07:14-0400

Solution:

The molar mass of H2O is 18.0 g mol-1.

Hence,

(24.93 g H2O) × (1 mol H2O / 18.0 g H2O) = 1.385 mol H2O

Moles of H2O = 1.385 mol


This problem can be summarized thusly:

1) Ice is heated from -16°C to 0°C.

The heat absorbed is calculated by using the specific heat of ice (Sice = 2.1 J/g°C) and the equation:

q = m × S × ΔT

q1 = m × Sice × (Tf - Ti)

q1 = 24.93 g × 2.1 J/g°C × (0 - (-16))°C = 837.65 J = 0.83765 kJ

q1 = 0.83765 kJ


2) Ice is melted at 0°C. The heat absorbed is calculated by multiplying the moles of ice by the molar heat of fusion (ΔHf).

q2 = ΔHf × n = 6.02 kJ/mol × 1.385 mol = 8.3377 kJ

q2 = 8.3377 kJ


3) Water at 0°C is heated to 100°C. The heat absorbed is calculated by using the specific heat of water (Swater = 4.2 J/g°C) and the equation: q = m × S × ΔT

q3 = m × Swater × (Tf - Ti)

q3 = 24.93 g × 4.2 J/g°C × (100 - 0)°C = 10470.6 J = 10.4706 kJ

q3 = 10.4706 kJ


4) Water is vaporized to steam at 100°C. The heat absorbed is calculated by multiplying the grams of water by the heat of vaporization (ΔHv) in kJ/g.

q4 = ΔHv × m = 2.26 kJ/g × 24.93 g = 56.3418 kJ

q4 = 56.3418 kJ


5) Steam is heated from 100°C to 113°C. The heat absorbed is calculated by using the specific heat of steam (Ssteam = 1.7 J/g°C) and the equation: q = m × S × ΔT

q5 = m × Ssteam × (Tf - Ti)

q5 = 24.93 g × 1.7 J/g°C × (113 - 100)°C = 550.853 J = 0.55095 kJ

q5 = 0.55095 kJ


q = q+ q2 + q3 + q4 + q5

q = 0.83765 kJ + 8.3377 kJ + 10.4706 kJ + 56.3418 kJ + 0.55095 kJ = 76.5387 kJ = 76.54 kJ

q = 76.54 kJ


Answer: 76.54 kJ of energy was absorbed by the sample.